OpenStudy (hoopcoach99):

Your weight lifting class cost you a $20 fee up front and $5.50 per class after. The equation for this situation is y = 5.5x + 20. Make a T-table for the equation, using the x values 0, 10, and 20. Which ordered pairs are represented in your T-table? A. (0, 0), (10, 55), (20, 110) B. (0, 0), (55, 10), (110, 20) C. (0, 20), (10, 75), (20, 130) D. (20, 0), (75, 10), (130, 20)

OpenStudy (hoopcoach99):

Can anyone help

OpenStudy (anonymous):

all you really have to do is plug these x values into the equation given

OpenStudy (hoopcoach99):

So just see if they add up

OpenStudy (anonymous):

yea.. the answer you get would be your y coordinate

OpenStudy (hoopcoach99):

So if I did it correct I got A as my answer

OpenStudy (hoopcoach99):

is that right

OpenStudy (anonymous):

not quite

OpenStudy (hoopcoach99):

Oh, do I have to add the 20 also

OpenStudy (anonymous):

yup

OpenStudy (hoopcoach99):

so C

OpenStudy (anonymous):

yup^^

OpenStudy (hoopcoach99):

Thank you, can you show me how to do this one also-Write the equation y - 2 = 4(x + 5) in standard form. A. 4x + y = 7 B. 4x - y = 3 C. 4x - y = -22

OpenStudy (anonymous):

standard form of a line =Ax+By=C

OpenStudy (anonymous):

take the first equation you have and simplify it

OpenStudy (hoopcoach99):

so take 4x+By=7

OpenStudy (anonymous):

not quite.. take y-2=4(x+5) and simplify you would get?

OpenStudy (hoopcoach99):

y-2=4x+20

OpenStudy (anonymous):

alright subtract the two on both sides

OpenStudy (hoopcoach99):

y=4x-18

OpenStudy (anonymous):

y=4x+22

OpenStudy (anonymous):

sorry i meant add..

OpenStudy (hoopcoach99):

ok, so 4x-y=22

OpenStudy (anonymous):

-22

OpenStudy (hoopcoach99):

yep, I forgot the -

OpenStudy (hoopcoach99):

thank you

OpenStudy (anonymous):

np